Tia, Zarko, Xiao, Tom, Mike, Bill, Cory, Zarko, Sam, Carrie, Greg

Horn 4 Target Systems: Tom on shift, will be joining from now on.

Booster: Finishing up shutdown work. Think it will come up fast, 4e12 dampers need to become commissioned. Beam early next week. Working on absorbers. Several days of studies. Regular ramping to Booster, check out. After comfortable, will go to 2-3 days of DC beam settings (stay at injection energy). Tom would like to do trim study at low intensity. Uninterrupted during owl shifts normally. 

BNB: Waiting for MI LCW. Raw skid fed from external chiller. May take ~1week to finish the system to run the magnets. Run through with safety today. Final sign off on Friday. Close up caverns. LCW should be ready Monday to do trim magnet tests.

Horn4: Bob, exterior half shells being rolled at vendor. Looking at end of Nov. for completion. Organize the test pulses for this with MicroBooNE install top panels of Cosmic Tagger System. As long as NuMI shutdown is finished, plenty of resources to do this test in Nov. Half time from pulse test, find issues that need to be fixed. (MI will not come up to 700 as fast as people think, probably by February, consider how this may impact testing of Horn 4 and coordinating with uB.)

Power outage: Jan/Feb switch Kautz rd power supply.

g-2: invite them to this meeting. Will be taking some beam in spring, mostly after next shutdown.

ANNIE: Bearing went bad in sump pump. Off for now, waiting for new one. Doing water quality tests. Installing cosmic tagger.

MicroBooNE: Wrapping up shutdown activities. Test PMT powersupply. Oct. 5-10, detector checkout dates. Maybe pit-level CRT work. Will be ready for beam on the tenth. Keep updated on high intensity beam for timing tests.